### v1

@@ 2 Methodology
Moreover, we observe that the delay in the step-response is significant and is an effect of the low-pass behavior indicated by the motors. This behavior greatly affects the dynamics of the Crazyflie nano quadrotor used in this study. According to the manufacturer’s measurements [^2] Crazyflie motor step response: https://www.bitcraze.io/2015/02/measuring-propeller-rpm-part-3/, we have determined through empirical analysis that a time interval of $0.15$ seconds would be sufficient for sim2real transference. These delays are considerably longer than the typical control interval of low-level controllers in Crazyflie. Therefore, they result in actions that affect the state only after $5$ to $25$ control steps. In order to tackle the major amount of partial observability, we include the history of control actions into the observation. This is a proprioceptive measurement that can be easily done in software without the need of additional hardware. In addition, noise was introduced into the observation to imitate imperfections in the sensor feedback of the reality platform. The unexpected noise values have a standard deviation of $0.001$ for position and orientation, and $0.002$ for linear velocity and angular velocity, respectively.
The observations are $18+4\cdot N_{H_{a}}$ dimensional, where $N_{H_{a}}$ is the number of action history, described as follows:
@@ 4 Conclusion
ACKNOWLEDGEMENT
This work was supported by Future Space Navigation & Satellite Research Center through the National Research Foundation funded by the Ministry of Science and ICT, the Republic of Korea (2022M1A3C2074404).

### v2

@@ 2 Methodology
Moreover, it is observed that the delay in the step-response is significant and is an effect of the low-pass behavior indicated by the motors. This behavior greatly affects the dynamics of the Crazyflie nano quadrotor used in this study. According to the manufacturer’s measurements [^2] Crazyflie motor step response: https://www.bitcraze.io/2015/02/measuring-propeller-rpm-part-3/, we have determined through empirical analysis that a time interval of $0.15$ seconds would be sufficient for sim2real transference. These delays are considerably longer than the typical control interval of low-level controllers in Crazyflie. Therefore, they result in actions that affect the state only after $5$ to $25$ control steps [18]. In order to tackle the major amount of partial observability, we include the history of control actions into the observation. This is a proprioceptive measurement that can be easily done in software without the need of additional hardware. In addition, noise was introduced into the observation to imitate imperfections in the sensor feedback of the reality platform [18]. The unexpected noise values have a standard deviation of $0.001$ for position and orientation, and $0.002$ for linear velocity and angular velocity, respectively.
The observations are $18+4\cdot N_{H_{a}}$ dimensional, where $N_{H_{a}}$ is the number of action history [18], described as follows: